OpenStudy (anonymous):

What is the exact area of a circle whose diameter is 24 cm?

OpenStudy (anonymous):

the area of a circle is A=πr^2

OpenStudy (anonymous):

150.72?

OpenStudy (anonymous):

\(\large\color{#c910a6}{A=πr^2}\) the radius is half of the diameter so r=12. \(\large\color{#c910a6}{A=π12^2}\)

OpenStudy (anonymous):

nope. what is 12 * 12 ?

OpenStudy (anonymous):

144

OpenStudy (dahlindubs11):

3.14159265358979x144=452.38934211692976

OpenStudy (anonymous):

yep times pi. (use calculator)

OpenStudy (anonymous):

452.16

OpenStudy (anonymous):

but thats not a answer

OpenStudy (anonymous):

452.39 but close enough.

TheSmartOne (thesmartone):

It wants the exact area, so just leave it with pi

TheSmartOne (thesmartone):

\(\sf\large 144\pi\) is the exact value. Not some decimal number.

OpenStudy (anonymous):

A. 576cm B.144 C.48 D.24

TheSmartOne (thesmartone):

Copy pasting?

OpenStudy (anonymous):

huh?

TheSmartOne (thesmartone):

When you copy paste, you do realize that pi is an image in your answer choice and that you can not copy paste images?

OpenStudy (anonymous):

yes.

TheSmartOne (thesmartone):

Please do not copy and paste your problems. If people are going to spend time to help you, you should at least spend the time to ask your question as clearly as possible. ~ http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok